OpenStudy (yanasidlinskiy):

Statistics? Anyone? I'd appreciate the help please:) The grade appeal process at a university requires that a jury be structured by selecting seven individuals randomly from a pool of seven students and seven faculty.​ (a) What is the probability of selecting a jury of all​ students? (b) What is the probability of selecting a jury of all​ faculty? (c) What is the probability of selecting a jury of four students and three ​faculty?

OpenStudy (zarkon):

do you know the hypergeometric distribution

OpenStudy (yanasidlinskiy):

Like the formula?

OpenStudy (zarkon):

it is a probability mass function

OpenStudy (yanasidlinskiy):

Ok, then yes I have.

OpenStudy (zarkon):

you can use that to answer the question

OpenStudy (yanasidlinskiy):

The problem is, is that I don't know how to use that and everything is just confusing me. It's my first attempt so I don't know.

OpenStudy (zarkon):

do you know this notation \[{n\choose r}\]

OpenStudy (yanasidlinskiy):

yea.

OpenStudy (zarkon):

ok P(a students and 7-a faculty chosen) \[=\Large\frac{{7\choose a}{7\choose 7-a}}{{14\choose 7}}\]

OpenStudy (yanasidlinskiy):

Okay, so what's my next step?

OpenStudy (zarkon):

replace a with the numbers you need

OpenStudy (yanasidlinskiy):

Where do I get the numbers from.... I only have 7 and 14. Unless I'm supposed to add them 2.

OpenStudy (zarkon):

"What is the probability of selecting a jury of all​ students? " all students means a=7 and \[=\Large\frac{{7\choose 7}{7\choose 7-7}}{{14\choose 7}}\]

OpenStudy (zarkon):

\[=\Large\frac{{7\choose 7}{7\choose 0}}{{14\choose 7}}\]

OpenStudy (yanasidlinskiy):

What would I do then?

OpenStudy (zarkon):

I've given you one of the answers (after you calculate it from my formula) I'm not giving all the answers

OpenStudy (yanasidlinskiy):

No, I know. I'm just asking because that formula really confuses me. I've been given this:\[_{n}C _{r} = \frac{ n! }{ r!(n-r)! }\]

OpenStudy (zarkon):

\[{n\choose r}= ~_{n}C_r\]

OpenStudy (zarkon):

same thing

OpenStudy (yanasidlinskiy):

Ohh, wow. Makes sense now. Thanks.

OpenStudy (zarkon):

\[=\Large\frac{{7\choose 7}{7\choose 0}}{{14\choose 7}}\] \[=\Large\frac{_7C_7\cdot _7C_0}{{_{14}C_7}}\]

OpenStudy (yanasidlinskiy):

Got it! That's what I needed thanks.